OpenStudy (anonymous):

solve You have $59.95 in your wallet and want to buy some new CDs. If the CDs are $11.99 each, what number of CDs, x, can you buy?

OpenStudy (anonymous):

try 11.99(5)

OpenStudy (anonymous):

that would equal $59.95 so you can buy 5 cds.

OpenStudy (anonymous):

I need to show step by step. So would it be x over 11.99 = 59.95 divided by 11.99

OpenStudy (anonymous):

If you were to do it that way, it wouldn't equal $59.95, would it? ...but you could do 59.95 over 11.99 and that would equal 5 which is what x is. Try dividing 11.99 by both sides 11.99/11.99x=59.95/11.99 and that gives you x=5

OpenStudy (anonymous):

got it thanks some much

OpenStudy (anonymous):

No problem(: